# Memory-Efficient Sequential Pattern Mining with Hybrid Tries

### v2

@@ Abstract
As modern data sets continue to grow exponentially in size, the demand for efficient mining algorithms capable of handling such large data sets becomes increasingly imperative. This paper develops a memory-efficient approach for Sequential Pattern Mining (SPM), a fundamental topic in knowledge discovery that faces a well-known memory bottleneck for large data sets. Our methodology involves a novel hybrid trie data structure that exploits recurring patterns to compactly store the data set in memory; and a corresponding mining algorithm designed to effectively extract patterns from this compact representation. Numerical results on real-life test instances show an average improvement of 88% in memory consumption and 41% in computation time for small to medium-sized data sets compared to the state of the art. Furthermore, our algorithm stands out as the only capable SPM approach for large data sets within 256GB of system memory.
Keywords Sequential pattern mining, Memory efficiency, Large-scale pattern mining, Trie data set models.
@@ 1 Introduction
Data volume is growing at an exponential rate (Taylor, 2021), with modern machine learning data sets often containing trillions of data points (Villalobos and Ho, 2022). While supervised machine learning algorithms have thrived by training on such large data sets, unsupervised algorithms face ongoing challenges in scalability due to their memory requirements. In particular, Sequential Pattern Mining (SPM), a prominent topic in unsupervised learning, encounters a well-known memory bottleneck in its two most prevalent algorithms. The Apriori algorithm (Agrawal et al., 1994) suffers from the explosion of candidate patterns, and the prefix-projection algorithm (Han et al., 2001) requires fitting the entire training data set into memory. This has limited extant SPM algorithms to smaller-sized data sets and impractical for larger data sets (Pillai and Vyas, 2011).
@@
The primary approach to applying SPM algorithms on large data sets is by using more hardware, either on a single computing unit or via a parallelized structure (such as the works of Gan et al. (2019); Huynh et al. (2018); Chen et al. (2017); Yu et al. (2019); Saleti and Subramanyam (2019); Chen et al. (2013)). However, using additional hardware is costly, requires specialized machinery in the case of parallelization, and is often capped due to technology and other system specifications. Maximizing performance under capped memory is thus a practical use-case that is of interest in SPM (Wang et al., 2003), and has been shown to be beneficial in various machine learning techniques (Pleiss et al., 2017; Gruslys et al., 2016; Si et al., 2017). Benefits include, for example, faster data set access and higher time efficiency, reduced overhead and optimized resource utilization, reduced hardware maintenance costs, facilitation of use in dynamic and online data streams, reduced environmental impact, and higher energy efficiency. Most importantly, a memory-efficient algorithm makes SPM available to a broader range of users who are constrained by hardware limitations.
@@
Motivated by the benefits mentioned above, this paper aims to enhance the memory efficiency of SPM algorithms, while preserving their time efficiency. To this end, we focus on the prefix-projection algorithm, which has been shown to improve over other alternative SPM algorithms (Han et al., 2001) in terms of time efficiency. The memory bottleneck of prefix projection arises from the necessity to fit the entire data set into memory, prompting a critical examination of how sequential data sets are modeled and stored. The dominant approach is to model the data set using a relational or vector model, such as the ones discussed in Fournier-Viger et al. (2017). An advantage of vector models is their simple structure, leading to a straightforward mining process. Nevertheless, a major disadvantage of vector models is their high memory usage which becomes increasingly inefficient as data sets grow in size. This has led researchers to explore alternative models such as trie structures, known for their concise representation of strings.
@@
A trie is a graph-based data structure commonly used to store associative arrays or sets of strings in an efficient manner. Trie models of data sets have been shown to provide benefits in mining algorithms, such as faster item-set mining (Han et al., 2004; Pyun et al., 2014; Borah and Nath, 2018), effective Apriori and candidate pattern storage (Ivancsy and Vajk, 2005; Pyun et al., 2014; Huang et al., 2008; Fumarola et al., 2016; Antunes and Oliveira, 2004; Wang et al., 2006; Bodon and Rónyai, 2003; Masseglia et al., 2000), effective web access mining (Yang et al., 2007; Pei et al., 2000; Lu and Ezeife, 2003), mining long biological sequences (Liao and Chen, 2014), up-down SPM (Chen, 2009), incorporating constraints into SPM (Masseglia et al., 2009; Hosseininasab et al., 2019), progressive SPM (Huang et al., 2006; El-Sayed et al., 2004), and faster SPM (Rizvee et al., 2020).
A major advantage of trie models is their ability to model multiple overlapping subsequences using only a single path. This ability is essential to modeling larger data sets and has been shown to achieve higher time and memory efficiency in the mining algorithm (Mabroukeh and Ezeife, 2010). Conversely, when only a few subsequences overlap, for example, in smaller data sets or ones with longer sequences, the memory overhead of modeling a sequence by a trie path may be higher than its vector representation.
@@
While trie models can provide a more memory-efficient model of the data set, using that model for SPM poses an entirely different challenge. Unlike vector models which have a one-to-one correspondence between their vectors and sequences of the data set, a single path of a trie model may correspond to several sequences of the data set. This makes tracking the frequency of patterns nontrivial, requiring additional information to be stored at the nodes of the trie. In fact, what information is kept at the nodes of a trie model and how it is used in the mining algorithm is the distinguishing factor between the many trie-based approaches in the literature. Unfortunately, all current trie models involve storing a rich set of information at their nodes in favor of time efficiency, increasing their memory requirements often beyond that of vector models.
@@
For example, in web access mining (a special case of SPM with simplified sequence structures), Pei et al. (2000) propose to use a hash table of linked nodes to traverse and scan their trie data set model. Using hash tables, the data set is recursively projected onto conditional smaller tries and mined accordingly. The disadvantage of this approach is the memory overhead of the hash table and the memory and time spent constructing the conditional tries. Instead of constructing conditional tries, Lu and Ezeife (2003) propose to store at each node of the trie model an integer position value and Yang et al. (2007) propose to recursively generate sub-header tables in the mining algorithm. Such integer position values grow exponentially with the size of the trie, and similarly, the generation of many subheader tables leads to higher memory consumption.
@@
Following works in web access mining, Rizvee et al. (2020) extend the trie model of a sequential data set to accommodate SPM. This is a challenging task, due to the different structural properties of data sets in SPM compared to web access mining (Mabroukeh and Ezeife, 2010). Their TreeMiner algorithm uses a similar idea to hash tables, and stores at each node of the trie a matrix of links that are used to traverse and mine the trie. Although this improves the time efficiency of the TreeMiner algorithm, the matrix of links grows linearly with the size of the data set, making the algorithm costly in memory usage.
@@
To improve over the high memory usage of vector and trie models, we begin by introducing a new binary trie model of the data set that stores a constant amount of information at each node. We prove that this information suffices for the subsequent mining task, and develop a novel algorithm with significantly lower memory requirement than that of Rizvee et al. (2020). Furthermore, we prove that our trie model is always asymptotically smaller than a vector model, giving it a major advantage in SPM of large data sets.
@@
In the next step, we build on our binary trie model and develop a hybrid trie-vector model of the data set. The idea is to take advantage of the strengths of both trie and vector models and further increase the time and memory efficiency of the subsequent SPM task. In particular, we exploit the fact that data set sequences have high overlap in their initial entries (that is, their prefixes), which can be effectively modeled via a trie, and low overlap in their latter entries (that is, their suffixes), which are more efficiently modeled via a vector. Our hybrid model is thus designed to find an effective transition from a trie model of a prefix to a vector model of the suffix when the compression provided by a trie representation does not compensate for its memory overhead.
@@
Lastly, we integrate our trie-mining algorithm and the prefix-projection algorithm into a novel hybrid mining algorithm that can accommodate and mine our hybrid data set structure. We experimentally show that our hybrid algorithm outperforms both trie and vector models in time and memory usage. In particular, our hybrid algorithm can model and mine orders of magnitude larger data sets that are out of reach for any other SPM algorithm. Although primarily designed to handle large data sets, our algorithm also shows improvements in time and memory usage when applied to small to medium-sized data sets.
@@ 2 Preliminaries
This section provides preliminaries on SPM, starting with the definition of vector models of sequential data sets and followed by the definition of trie models. Throughout the paper, and for a thorough space complexity analysis, we consider that an integer $j$ has $\mathcal{O}(\log(|j|))$ space complexity (Papadimitriou and Steiglitz, 1998). We also discuss the relaxation of this assumption for reasonably-sized integers that can be stored in memory using constant overhead.
@@ 2.1 Vector Models of the Data Set for SPM
Let $\mathbb{E}=\left\{e_{1},\dots,e_{|\mathbb{E}|}\right\}$ be a finite set of literals, representing possible events or items within an application of interest. An itemset $\mathbb{I}=\left\{e_{1},\dots,e_{\left|\mathbb{I}\right|}\right\}$ is a set of events such that $\mathbb{I}\subseteq\mathbb{E}$. Although events of an itemset can be of any order, they are generally assumed to satisfy a monotone property (Han et al., 2001) and ordered accordingly. A sequence $\mathcal{S}_{i}=\left\langle\mathbb{I}_{i}^{1},\dots,\mathbb{I}_{i}^{L_{i}}\right\rangle$ is an ordered list of $L_{i}$ itemsets, with size $|\mathcal{S}_{i}|=\sum_{j=1}^{L_{i}}|\mathbb{I}_{i}^{j}|$. An event $e\in\mathbb{E}$ can occur at most once in an itemset, but the same itemset can occur multiple times in a sequence. Sequences may be equivalently represented in the event space $\mathcal{S}_{i}=\left\langle\bar{e}_{1},\dots,e_{j},\bar{e}_{j+1},\dots,e_{|\mathcal{S}_{i}|}\right\rangle$ with the first event of any itemset indicated by an accented event $\bar{e}$.
A sequential data set $\mathcal{SD}$ is a list of $N$ sequences $\mathcal{SD}=\left\langle\mathcal{S}_{1},\dots,\mathcal{S}_{N}\right\rangle$, with the size of its largest sequence denoted by $M=\max_{i\in\{1,\dots,N\}}|\mathcal{S}_{i}|$. Example 1 describes a small instance of a sequential data set given in Table 1, which we use as a running example throughout this paper.
@@
The SPM task involves finding the set of frequent patterns within a data set $\mathcal{SD}$. A pattern $\mathcal{P}$ is a subsequence that satisfies $\mathcal{P}\sqsubseteq\mathcal{S}_{i}$ for at least one sequence $\mathcal{S}_{i}\in\mathcal{SD}$. The support $supp(\mathcal{P})\in\mathbb{Z}_{>0}$ of a pattern $\mathcal{P}$ is the number of distinct sequences for which $\mathcal{P}\sqsubseteq\mathcal{S}_{i}$. A pattern is considered frequent if $supp(\mathcal{P})\geq\theta\times N$, where $0<\theta\leq 1$ is a minimum support threshold.
Frequent patterns are generally found iteratively, where at each iteration a frequent pattern $\mathcal{P}$ is extended by a single event $e\in\mathbb{E}$ to $\langle\mathcal{P},e\rangle$ and checked to satisfy $supp(\langle\mathcal{P},e\rangle)\geq\theta\times N$. Pattern extensions are classified into an itemset extension or a sequence extension. In an itemset extension, pattern $\mathcal{P}$ is extended to $\langle\mathcal{P},e\rangle$, extending its last itemset $\mathbb{I}^{|\mathcal{P}|}$ by a single event. In a sequence extension, a pattern is extended to $\langle\mathcal{P},\bar{e}\rangle$, extending $\mathcal{P}$ by a new itemset $\mathbb{I}=\{\bar{e}\}$. Example 2 demonstrates a frequent pattern in our running example.
@@
The literature generally models sequential data sets via vectors, where each sequence $\mathcal{S}_{i}\in\mathcal{SD}$ is stored in memory using a single vector. Current state-of-the-art SPM algorithms that use vector models store the entire model in memory, which can be costly in terms of memory usage. In particular, Lemma 3 gives the worst-case space complexity of a vector-based SPM algorithm.
@@
A sequential data set $\mathcal{SD}$ uses one vector per sequence $\mathcal{S}_{i}\in\mathcal{SD}$ and stores an integer $j\leq|\mathbb{E}|$ per event $e\in\mathcal{S}_{i}$. The worst-case space complexity of a vector model is thus $\mathcal{O}\left(Nc^{\mathcal{V}}+NM\log(|\mathbb{E}|)\right)=\mathcal{O}\left(NM\log(|\mathbb{E}|)\right)$.
The most memory-efficient and basic SPM algorithm involves pseudo-projection (Han et al., 2001). Pseudo-projection involves storing a position pair $(i,j)$, where $i\leq N$ is a sequence ID and $j\leq M$ is a positional integer, for at most all $\mathcal{S}_{i}\in\mathcal{SD}$ and positions $M$. The worst-case space complexity of a vector-based SPM algorithm is thus $\mathcal{O}(NM\log(M)+NM\log(N))=\mathcal{O}(NM\log(N))$.
Therefore, the worst-case space complexity of vector-based SPM algorithms is $\mathcal{O}(NM\log(|\mathbb{E}|)+NM\log(N))=\mathcal{O}\left(NM\log(N)\right)$. Reasonably sized integers $N$ consume constant memory, and reduce the complexity to $\mathcal{O}\left(NM\right)$
As shown in Lemma 3, the memory consumption of vector models and algorithms grows linearly with $N$ for reasonably sized values of $N$. In practice, this can be considerably memory-consuming, for example, for the data sets of Villalobos and Ho (2022); Criteo AI Lab (2014); Consortium et al. (2015) that include billions to trillions of sequences. Consequently, vector-based SPM algorithms cannot often fit large datasets into system memory, preventing them from performing their subsequent mining task. To improve on this memory deficiency, we next examine trie models, which have the potential to conserve memory by representing multiple sequences of the sequential data set using a single path.

### v3

@@ Abstract
This paper develops a memory-efficient approach for Sequential Pattern Mining (SPM), a fundamental topic in knowledge discovery that faces a well-known memory bottleneck for large data sets. Our methodology involves a novel hybrid trie data structure that exploits recurring patterns to compactly store the data set in memory; and a corresponding mining algorithm designed to effectively extract patterns from this compact representation. Numerical results on small to medium-sized real-life test instances show an average improvement of 85% in memory consumption and 49% in computation time compared to the state of the art. For large data sets, our algorithm stands out as the only capable SPM approach within 256GB of system memory, potentially saving 1.7TB in memory consumption.
Keywords :
Sequential pattern mining, Memory efficiency, Large-scale pattern mining, Trie data set models.
@@ 1 Introduction
Data volume is growing at an exponential rate (Taylor, 2021), with modern machine learning data sets often containing trillions of data points (Villalobos and Ho, 2022). While supervised machine learning algorithms have thrived by training on such large data sets, unsupervised algorithms face ongoing challenges in scalability due to their memory requirements. In particular, Sequential Pattern Mining (SPM), a prominent topic in unsupervised learning, encounters a well-known memory bottleneck in its two most prevalent algorithms. The Apriori algorithm (Agrawal et al., 1994) suffers from the explosion of candidate patterns that are costly to store in memory, and the prefix-projection algorithm (Han et al., 2001) requires fitting the entire training data set into memory. This has limited extant SPM algorithms to smaller-sized data sets and rendered them impractical for larger ones (Pillai and Vyas, 2011).
@@
The primary approach to applying SPM algorithms on large data sets is by using more hardware, either on a single computing unit or via a parallelized structure, such as Gan et al. (2019); Huynh et al. (2018); Chen et al. (2017); Yu et al. (2019); Saleti and Subramanyam (2019); Chen et al. (2013). However, using additional hardware is costly, requires specialized machinery in the case of parallelization, and is often capped due to technology and other system specifications. Maximizing performance under capped memory is thus a practical use-case that is of interest in SPM (Wang et al., 2003), and has been shown to be beneficial in various machine learning techniques such as Pleiss et al. (2017); Gruslys et al. (2016); Si et al. (2017). Benefits include, for example, faster data set access and higher time efficiency, reduced overhead and optimized resource utilization, reduced hardware maintenance costs, facilitation of use in dynamic and online data streams, reduced environmental impact, and higher energy efficiency. Most importantly, a memory-efficient algorithm makes SPM available to a broader range of users who are constrained by hardware limitations.
@@
Motivated by these benefits, this paper aims to enhance the memory efficiency of SPM algorithms, while simultaneously preserving or increasing their time efficiency. We focus on the prefix-projection algorithm, which has been shown to improve over other alternative SPM algorithms in terms of time efficiency (Han et al., 2001). The memory bottleneck of prefix-projection arises from the necessity to fit the entire data set into memory, prompting a critical examination of how sequential data sets are modeled and stored. The dominant approach is to model the data set using a relational or vector model, such as the ones discussed in Fournier-Viger et al. (2017). An advantage of vector models is their simple structure, leading to a straightforward mining process. Nevertheless, a major disadvantage of vector models is their high memory usage which becomes increasingly inefficient as data sets grow in size. This has led researchers to explore alternative models such as trie structures which are known for their concise representation of strings.
@@
A trie is a graph-based data structure commonly used to store associative arrays or sets of strings in an efficient manner. A major advantage of trie models is their ability to model multiple overlapping subsequences using only a single trie path, which can potentially lead to higher time and memory efficiency in the mining algorithm (Mabroukeh and Ezeife, 2010). For example, trie models of data sets have been shown to provide benefits such as faster item-set mining (Han et al., 2004; Pyun et al., 2014; Borah and Nath, 2018), effective Apriori and candidate pattern storage (Ivancsy and Vajk, 2005; Pyun et al., 2014; Huang et al., 2008; Fumarola et al., 2016; Antunes and Oliveira, 2004; Wang et al., 2006; Bodon and Rónyai, 2003; Masseglia et al., 2000), effective web access mining (Yang et al., 2007; Pei et al., 2000; Lu and Ezeife, 2003), mining long biological sequences (Liao and Chen, 2014), up-down SPM (Chen, 2009), incorporating constraints into SPM (Masseglia et al., 2009; Hosseininasab et al., 2019; Wang et al., 2022; Kadıoğlu et al., 2023), progressive SPM (Huang et al., 2006; El-Sayed et al., 2004), and faster SPM (Rizvee et al., 2020).
@@
While trie models can provide a more memory-efficient model of the data set, using that model for SPM poses an entirely different challenge. Unlike vector models which have a one-to-one correspondence between their vectors and sequences of the data set, a single path of a trie model may correspond to several sequences. This makes tracking the frequency of patterns nontrivial, requiring additional information to be stored at the nodes of the trie. In fact, the distinguishing factor between the many trie-based approaches in the literature is the information stored at the nodes of their tries and how it is used in the mining algorithm. Unfortunately, all current trie models involve storing a rich set of information at their nodes in favor of time efficiency, often increasing their memory requirements beyond that of vector models.
@@
For example, in web access mining (a special case of SPM with simplified sequence structures), Pei et al. (2000) propose to use a hash table of linked nodes to traverse and scan their trie data set model. Using hash tables, the data set is recursively projected onto conditional smaller tries and mined accordingly. The disadvantage of this approach is the memory overhead of the hash table and the memory and time spent constructing the conditional tries. Instead of constructing conditional tries, Lu and Ezeife (2003) propose to store at each node of the trie model an integer position value, and Yang et al. (2007) propose to recursively generate sub-header tables in the mining algorithm. Such integer position values grow exponentially with the size of the trie, and similarly, the generation of many sub-header tables leads to higher memory consumption.
@@
Following works in web access mining, Rizvee et al. (2020) extend the trie model of a sequential data set to accommodate SPM. This is a challenging task, due to the different structural properties of data sets in SPM compared to web access mining (Mabroukeh and Ezeife, 2010). Their TreeMiner algorithm uses a similar idea to hash tables, and stores at each node of the trie a matrix of links that are used to traverse and mine the trie. Although this can improve the time efficiency of the TreeMiner algorithm, the matrix of links grows linearly with the size of the data set and makes the algorithm costly in memory usage.
@@
To improve over the high memory usage of vector and trie models, we begin by introducing a new binary trie model of the data set that stores a constant amount of information at each node. We prove that this information suffices for SPM, and develop a novel algorithm with significantly lower memory requirement than that of Rizvee et al. (2020). Furthermore, we prove that our trie model is always asymptotically smaller than a vector model, giving it a major advantage in SPM of large data sets. On the other hand, in data sets where only a few subsequences overlap, such as smaller data sets or ones with longer sequences, the memory overhead of modeling a sequence by a trie path may be higher than its vector representation.
@@
To improve memory efficiency for such data sets, we build on our binary trie and develop a hybrid trie-vector model of the data set. The idea is to take advantage of the strengths of both trie and vector models and further increase the time and memory efficiency of the subsequent SPM task. In particular, we exploit the fact that data set sequences have high overlap in their initial entries (that is, their prefixes)—which can be effectively modeled via a trie—and low overlap in their latter entries (that is, their suffixes)—which are more efficiently modeled via a vector. Our hybrid model is thus designed to find an effective transition from a trie model of prefixes to a vector model of suffixes that improves memory consumption.
@@
To mine our hybrid data set structure, we combine our binary trie mining algorithm with a vector-based prefix-projection algorithm to develop a novel hybrid mining algorithm. We experimentally show that our hybrid algorithm outperforms both trie and vector models in time and memory usage. In particular, our hybrid algorithm can model and mine orders of magnitude larger data sets that are out of reach for any other SPM algorithm. Although primarily designed to handle large data sets, our algorithm improves time and memory usage when applied to small to medium-sized data sets, showcasing its potential for larger data sets in time efficiency.
@@ 2 Preliminaries
This section provides preliminaries on SPM, starting with the definition of vector models of sequential data sets and followed by the definition of trie models. Throughout the paper, and for a thorough space complexity analysis, we consider that a vector of integers uses a constant $c^{\mathcal{V}}$ memory overhead, and that an integer $j$ has $\mathcal{O}(\log(|j|))$ space complexity (Papadimitriou and Steiglitz, 1998). We also discuss the relaxation of the later assumption for reasonably-sized integers that can be stored in memory using constant overhead.
@@ 2.1 Vector Models of the Data Set for SPM
Let $\mathbb{E}=\left\{e_{1},\dots,e_{|\mathbb{E}|}\right\}$ be a finite set of literals, representing possible events or items within an application of interest. An itemset $\mathbb{I}=\left\{e_{1},\dots,e_{\left|\mathbb{I}\right|}\right\}$ is a set of events such that $\mathbb{I}\subseteq\mathbb{E}$. Although events of an itemset can be of any order, they are generally assumed to satisfy a monotone property and ordered accordingly (Han et al., 2001). A sequence $\mathcal{S}_{i}=\left\langle\mathbb{I}_{i}^{1},\dots,\mathbb{I}_{i}^{L_{i}}\right\rangle$ is an ordered list of $L_{i}$ itemsets, with size $|\mathcal{S}_{i}|=\sum_{j=1}^{L_{i}}|\mathbb{I}_{i}^{j}|$. An event $e\in\mathbb{E}$ can occur at most once in an itemset, but the same itemset can occur multiple times in a sequence. Sequences may be equivalently represented in the event space $\mathcal{S}_{i}=\left\langle\bar{e}_{1},\dots,e_{j},\bar{e}_{j+1},\dots,e_{|\mathcal{S}_{i}|}\right\rangle$, with the first event of any itemset indicated by an accented event $\bar{e}$. This notation is adopted throughout this paper.
A sequential data set $\mathcal{SD}$ is a list of $N$ sequences $\mathcal{SD}=\left\langle\mathcal{S}_{1},\dots,\mathcal{S}_{N}\right\rangle$, with the size of its largest sequence denoted by $M=\max_{i\in\{1,\dots,N\}}|\mathcal{S}_{i}|$. Example 1 describes a small instance of a sequential data set given in Table 1, which we use as a running example.
@@
The SPM task involves finding the set of frequent patterns within a data set $\mathcal{SD}$. A pattern $\mathcal{P}$ is a subsequence that satisfies $\mathcal{P}\sqsubseteq\mathcal{S}_{i}$ for at least one sequence $\mathcal{S}_{i}\in\mathcal{SD}$. The support $supp(\mathcal{P})\in\mathbb{Z}_{>0}$ of a pattern $\mathcal{P}$ is the number of distinct sequences in $\mathcal{SD}$ for which $\mathcal{P}\sqsubseteq\mathcal{S}_{i}$. A pattern is considered frequent if $supp(\mathcal{P})\geq\theta\times N$, where $0<\theta\leq 1$ is a user-defined minimum support threshold.
Frequent patterns are generally found iteratively, where at each iteration a frequent pattern $\mathcal{P}$ is extended by a single event $e\in\mathbb{E}$, to $\langle\mathcal{P},e\rangle$, and checked to satisfy $supp(\langle\mathcal{P},e\rangle)\geq\theta\times N$. Pattern extensions are classified into an itemset extension or a sequence extension. In an itemset extension, pattern $\mathcal{P}$ is extended to $\langle\mathcal{P},e\rangle$, extending its last itemset $\mathbb{I}^{|\mathcal{P}|}$ by a single event. In a sequence extension, a pattern is extended to $\langle\mathcal{P},\bar{e}\rangle$, extending $\mathcal{P}$ by a new itemset $\mathbb{I}=\{\bar{e}\}$. Example 2 demonstrates a frequent pattern in our running example.
@@
The literature generally models sequential data sets via vectors, where each sequence $\mathcal{S}_{i}\in\mathcal{SD}$ is stored in memory using a single vector. Current state-of-the-art SPM algorithms that use vector models store their entire representation of the data set in memory, which can be costly in terms of memory usage. In particular, Lemma 3 gives the worst-case space complexity of a vector-based SPM algorithm.
@@
A sequential data set $\mathcal{SD}$ uses one vector per sequence $\mathcal{S}_{i}\in\mathcal{SD}$ and stores an integer $j\leq|\mathbb{E}|$ per event $e\in\mathcal{S}_{i}$. Assuming that $|\mathbb{E}|\leq N$ and $M\leq N$ (which generally hold in practice), the worst-case space complexity of a vector model is $\mathcal{O}\left(Nc^{\mathcal{V}}+NM\log(|\mathbb{E}|)\right)=\mathcal{O}\left(NM\log(|\mathbb{E}|)\right)$.
The most memory-efficient and basic SPM algorithm involves pseudo-projection (Han et al., 2001). Pseudo-projection stores a sequence ID-integer position pair $(i,j)$ (with $i\leq N$ and $j\leq M$) for at most all $\mathcal{S}_{i}\in\mathcal{SD}$ and positions $M$. The worst-case space complexity of a vector-based SPM algorithm is thus $\mathcal{O}(NM\log(M)+NM\log(N))=\mathcal{O}(NM\log(N))$.
Putting the two together, the worst-case space complexity of vector-based SPM algorithms is $\mathcal{O}(NM\log(|\mathbb{E}|)+NM\log(N))=\mathcal{O}\left(NM\log(N)\right)$. Reasonably sized integers $N$ consume constant memory, and reduce the complexity to $\mathcal{O}\left(NM\right)$
As shown in Lemma 3, the memory consumption of vector models and algorithms grows linearly with $N$—for reasonably sized values of $N$. In practice, this can be highly memory-consuming, for example, for the large-size data sets of Villalobos and Ho (2022); Criteo AI Lab (2014); Consortium et al. (2015) that include billions to trillions of sequences. For such large data sets, vector-based SPM algorithms cannot fit their data set representation into system memory, which consequently prevents them from performing their mining task. To improve on this memory deficiency, we next examine trie models of the data set, which have the potential to conserve memory by representing multiple sequences of the sequential data set using a single path.